suppose the nominal rate of interest in an economy is 10% and rate of inflation is 6% per annum respectively,

a) determine the real rate of interest.

b) discuss two factors that lead to a change in the general level of interest rate.
1
Expert's answer
2014-09-23T12:02:17-0400
We can determine the real rate of interest using the Fisher equation that shows the relation between real and nominal interest rates and the expected inflation:
(1 + nominal rate) = (1 + real interest rate) (1 + inflation rate) So,
(1 + real interest rate) = (1 + nominal rate) / (1 + inflation rate) Real interest rate = (1 + nominal rate) / (1 + inflation rate) – 1 Real interest rate = (1 + 0.1) / (1 + 0.06) – 1 Real interest rate = 0.038 Real interest rate = 3.8% First of all, interest rates are influenced by the overall condition of the economy. When the economy is growing, consumers have jobs and savings to lend through banks. So, interest rates rise and act as a ration for the funds available due to the increases of demand for funds.
Also a great effect on interest rates has government. It can influence these rates using open market transactions. When it buys more securities, banks are injected with more money than they can use for lending, and the interest rates decrease and vice versa.